OpenStudy (anonymous):

What is the probability that the number cube will show a number greater than 4 and the coin will land with tails up?

OpenStudy (anonymous):

cube: 5 and 6 up is 2/6 probability coin: tail up is 1/2 probability both = (2/6)(1/2) = answer